OpenStudy (anonymous):

5/2t- t=3+3/2t??? Can someone help me???

OpenStudy (precal):

(5/2)t-(2/2)t=3+(3/2)t (3/2)t=3+(3/2)t 0=3 No solution No t value will make this true

OpenStudy (anonymous):

thanks so much :D